7m+n=30
m-3n=20
solve by elimination method

Found 2 solutions by mananth, Jk22:
Answer by mananth(16946)

7m+n=30
m-3n=20 Multiply by 7
7(m-3n)-(7m+n)=140-30
7m-21n-7m-n=110
-22n=110
n=-5
..
plug the value of n in the equation
m-3n=20
m+15=20
m=5

Answer by Jk22(389)

1) 7m+n=30
2) m-3n=20
2)=> m=20+3n
eliminate m in (1) : 7(20+3n)+n=140+21n+n=140+22n=30
=> 22n=-110 => n=-5
=> m=20-15=(30+5)/7=5